# Masked Particle Modeling on Sets: Towards Self-Supervised High Energy Physics Foundation Models

### v1

@@ I Introduction
While Artificial Intelligence (AI) and Machine Learning (ML) are already playing a major role in the analysis of high energy physics (HEP) data, the HEP community has yet to benefit from the self-supervised learning (SSL) based approaches to building large foundation models (FM) Bommasani et al. (2022) that have been pioneered in natural language processing (NLP) Lewis et al. (2019); Devlin et al. (2019); OpenAI (2023); Brown et al. (2020) and computer vision (CV) Dosovitskiy et al. (2021); Caron et al. (2021); Bao et al. (2022). These modern approaches use SSL to pre-train models on vast data sets in order to learn generic representations of the data. Such models can then be efficiently fine-tuned with small datasets for a variety of downstream tasks. The self-supervised pre-training of a FM produces a model that is also referred to as the “backbone", as it can serve as the information extraction component for downstream models. This concept significantly expands the possibilities for learning robust and meaningful data representations. The knowledge encoded in this way can be readily applied in downstream tasks which use this representation as input.
@@
The FM with a SSL pre-training approach offers important advantages for HEP. Unlike supervised learning, which typically acquires limited domain representations and focuses on a few key features for high prediction accuracy that must be learned anew for each task, SSL aims to learn generic representations summarizing domain features that prove useful across various downstream tasks. SSL tasks can be formulated on unlabeled data. In the HEP context, this may not only decrease the reliance on labeled simulated data sets but also potentially helps mitigate uncertainties related to domain shift when training models on imperfect simulations. However, this approach also has several major challenges for HEP. SSL strategies are data type specific, so new methods must be developed. These models also represent a scale in both model size and data size that have not been addressed in HEP. In this work, we aim to take the first steps towards building such a HEP foundation model, focusing on developing HEP data specific SSL strategies, whilst keeping an eye on how well such strategies may scale in the future. We propose a masked particle modeling (MPM) scheme, akin to masked language modeling (MLM) in NLP, for self-supervised learning on unlabeled data consisting of sets of particles in a collider physics environment. In doing so, we propose a novel scheme to apply masked modeling strategies to unordered sets of inputs.
Figure 1: The proposed model and training scheme for a FM for jets. A jet is represented as a set of particles, each a list of features, and some particles are replaced by masked tokens and passed through a transformer encoder. Training aims to predict the discrete token identity, defined by the encoder of a pre-trained VQ-VAE, of the masked particles.
@@
This work aims to generalize the language-inspired MLM-type training scheme to HEP scientific data. The paradigm involves extracting semantic meaning and understanding of the whole by predicting the missing (masked) pieces, referred to as tokens, thereby considering the collective impact of individual input elements. Generic data sets, such as the ones in HEP, cannot be directly mapped to a sequence of pieces in analogy to words in a sentence as they are continuous and unordered, i.e. not a sequence. The challenges of using continuous elements, i.e. features of particles like momentum, instead of a discrete dictionary of words, and of operating on unordered sets of inputs will be addressed in the development of the MPM scheme. To do so, we examine MPM in collider physics data sets consisting of jets, or collimated streams of particles produced by high-energy quarks and gluons.
@@ II Related Work
Foundation models, such as Masked Language Models (BART Lewis et al. (2019), BERT Devlin et al. (2019)), Generative Pre-trained Transformer (GPT) OpenAI (2023); Brown et al. (2020), Vision Transformer (ViT) Dosovitskiy et al. (2021), DINO Caron et al. (2021) and their combinations, such as DALLE Ramesh et al. (2021), Flamingo Alayrac et al. (2022) and others have primarily been explored in the domains of language and vision. We refer readers to the recent review Bommasani et al. (2022) for an overview. Most closely related to this work is the BERT model Devlin et al. (2019), which uses the masking and prediction of missing words as a pre-training task, and the BEiT model Bao et al. (2022), which adapts the masked language modeling method to images by masking and predicting patches of input images. On masked modeling schemes for data which consists of unorderd sets of inputs, the impact of removing positional information in masked image modeling was examined in Ref. Chen et al. (2021), and using position as a target when processing unordered image patches was explored in Ref. Zhai et al. (2022). The first steps in developing foundation models for science have been developed in e.g. protein biology Rives et al. (2021), molecular chemistry Ross et al. (2022); Pan (2023), and cosmology Lanusse et al. (2023); Walmsley et al. (2022), showing their ability to learn informative representations that are useful in these domains for various downstream tasks.
The first steps in self-supervised learning on jets was explored in Refs. Dillon et al. (2022, 2023); Tombs and Lester (2022), largely focusing on contrastive pre-training using augmentations of jets. Supervised pre-training strategies have also been explored in Ref. Kishimoto et al. (2023). Transformer models were trained on large jet datasets for classification in a supervised setting in Ref. Qu et al. (2022a); Mikuni and Canelli (2021) and several transformer-based applications have since been developed (for example, see Refs. K ach et al. (2022); Kansal et al. (2023); Fenton et al. (2022); ATLAS Collaboration (2023); Smith et al. (2023); Tomiya and Nagai (2023); K ach and Melzer-Pellmann (2023); Raine et al. (2023a)). Transformers have also been used for auto-regressive density estimation and jet classification Finke et al. (2023); Butter et al. (2023). Notably, Ref. Finke et al. (2023) also explored the discretization of continuous particle features to form jet sequences, which we examine in this work.
@@ III.1 Masked Particle Modeling
The representation of a jet as an unordered set of particles, each characterized by an unordered collection of continuous features, lends itself to interpretation within the masked training paradigm reminiscent of MLMs. In this analogy, each particle within the jet can be viewed as a representation akin to a token in a block of text, despite the notable difference that particles form unordered sets, while text is sequential.
@@
We can consider jets with at most $N$ constituent particles to be described as a set $X=\{x_{i}\}_{i=1}^{N}\in\mathcal{X}$ where each $x_{i}$ is some representation of a particle and $\mathcal{X}$ is the set of possible jets. A dataset is then a collection of $K$ jets $\{X^{j}\}_{j=1}^{K}$. The MPM objective can be phrased as partitioning the set of particles in each jet $j$ into a masked set $\mathcal{M}_{x}^{j}=\{x_{i}^{j}\}_{i\in\mathcal{M}^{j}}$ over indices $\mathcal{M}^{j}$ of masked elements, and an unmasked set $\mathcal{U}_{x}^{j}=\{x_{i}^{j}\}_{i\in\mathcal{U}^{j}}$ over indices of non-masked elements. A masking strategy is used for mapping $\mathcal{M}_{x}^{j}\rightarrow\mathcal{M}_{m}^{j}=\{m_{i}^{j}\}_{i\in\mathcal{M}^{j}}$ where each $m_{i}^{j}$ is referred to as a masked token. For convenience, we drop the jet index $j$ as the procedure is repeated for all jets. The goal is to define a parametric function $f_{\theta}:\mathcal{X}\to\mathbb{R}^{N\times d}$, which maps a jet to a $d$-dimensional representation for each of the $N$ particles in the jet, and a loss $\mathcal{L}$ such that minimizing $\mathbb{E}_{\mathcal{X}}[\mathcal{L}(\mathcal{M}_{x},f_{\theta}(\mathcal{M}_{m},\mathcal{U}_{x}))]$ results in a function $f_{\theta}$ that is useful for downstream tasks. As the goal of the pre-training is to recover information about each masked particle, we use a per-particle loss function of the form
@@
The particles that form a jet are permutation invariant, and so it is natural to use a backbone model that is permutation equivariant. However, if the same masked token is used to replace all of the particles in the masked set, then the output of the model will be equivalent at every masked position. In Fig. 1 this would mean that if $m_{1}=m_{2}$ then $h_{3}=h_{5}$, because permutation equivariant functions can not distinguish between two input particles with the same values. This exposes a redundancy that can only be removed by inducing an ordering, or by adding attributed connections between particles. The redundancy makes it impossible to achieve zero error on any single prediction unless the original value of every particle in the masked set is identical. However, the redundancy does not make MPM fruitless, as learning the density over the masked set of particles is still a useful and difficult task.
@@ III.2 Making Tokens From Particles
A scheme for defining context dependent tokens from continuous inputs was developed in the masked image modeling approach of the BEiT model Bao et al. (2022). In this case, labels of different image patches were defined using a Vector Quantized Variational AutoEncoder (VQ-VAE) Oord et al. (2017). A VQ-VAE uses an encoder to map a set of inputs to latent vectors, which are subsequently projected onto the nearest element within a finite codebook. These codebook vectors are then decoded back to the original inputs. The codebook vectors are trained simultaneously with the encoder and decoder. The use of transformer encoder and decoder ensures that information from all input elements are used to define the predictions. This process incorporates input-wide context into the codebook prediction for each element within the input set. In MPM, each particle is encoded to a single codebook element, where the encoding is performed conditionally on all other particles in the jet. The index of the codebook element to which each particle is mapped is then used as the target label during pre-training.
We also explore using the K-means algorithm MacQueen et al. (1967) to define labels for the pre-training. We use the K-means++ Arthur and Vassilvitskii (2007) algorithm as implemented in the scikit learn library Buitinck et al. (2013) to define the clusters. After training, each cluster is assigned an index, and the index is used as target labels. This approach is explored as a replacement for the VQ-VAE defined labels.
@@ V Experiments
The same transformer encoder architecture is used for all backbone models. We used transformer blocks based on the Normformer Shleifer et al. (2021). Eight transformer blocks are used with a model dimension of $1024$, the input nodes are embedded linearly into a $512$ dimensional space, and the model contains a total of 40 million parameters. The pre-training head is a single transformer block followed by a softmax for classification or a linear layer for regression. The fine-tuning classification head is a weighted average over all output dimensions (i.e. a linear transform applied independently to each particle representation and then averaged) followed by another linear layer and softmax. Models are trained with AdamW Kingma and Ba (2014); Loshchilov and Hutter (2017) with a learning rate of $10^{-4}$ and weight decay of $0.01$. All pre-trained models are trained for five epochs on the full JetClass data set. For supervised training, all models are trained for up to $50$ epochs with early stopping and a patience of five epochs. The norm of the gradient vectors are clipped to five. The from scratch supervised models use the same architecture as the pre-trained models: eight transformer blocks followed by a weighted average and softmax for classification. More details can be found in App. A.
@@
To define target labels and for input quantization, 512 possible token values are used. This corresponds to 512 vectors in the VQ-VAE codebook, or to 512 clusters for the kNN.
@@ V.1 Token Creation with a VQ-VAE
One challenge with training VQ-VAEs is the possibility of codebook collapse, where only a few of the codebook vectors are used to encode particle features. In the case of collapse to a single vector, the model would effectively be learning an average representation, which would not aid pre-training. To address this the prescriptions outlined in Ref. Huh et al. (2023) are applied as described in App. B. After training we find that the codebook has $\sim 80\%$ utilization.
@@ V.2 Discretization and Ordering
To study the impact of ordering, input quantization, and output quantization / loss function choice, backbone models are pre-trained under several different settings. The possible ordering settings included are: not ordering anywhere in the model, ordering the input to the backbone, and ordering only at the input to the pre-training prediction head. Ordering is defined by decreasing particle transverse momentum, and the index within the ordered sequence is embedded using learned positional embeddings Gehring et al. (2017); Vaswani et al. (2017). Learned positional embedding assigns a learnable vector to each index value that is then added to the features (and thus has the same dimension as the features per element). Depending on the model under study, a learned positional embedding is added to the backbone input features or the prediction head input features. Continuous inputs are used in all models except one, which tests the impact of quantizing particle feature inputs. For this input-quantized case, the codebook vectors assigned by the VQ-VAE are used to replace the input vectors from the data set. Finally, the classification loss using VQ-VAE encoded tokens as targets, the classification loss using kNN encoded tokens as targets, and the direct particle feature regression are tested.
@@
The results of the comparison can be seen in Tab. 1 where five different random seeds were used during the fine tuning stage, but with only a single backbone trained in each case. The mean result is shown in the table, with a standard deviation of O(0.01%) for each result . A clear benefit to ordering only in the model head is observed over ordering the backbone inputs or not ordering anywhere in the model. In this case, ordering the prediction head enables the symmetry over masked elements to be broken and thus make more precise per-particle predictions rather than a joint posterior over all masked particles. Input quantization is observed to hurt model performance, owing to the loss of information when summarizing continuous features with discrete tokens. Using tokens, either from the VQ-VAE or the kNN, as classification targets is found to be substantially better than regressing particle features as a prediction task. This is likely due to the regression giving only a point prediction of average particle properties under the posterior, which provides little predictive power for posteriors with multiple modes. The classification loss is much more flexible and able to make use of the prediction of the full posterior over tokens. The VQ-VAE tokens outperformed the kNN tokens for classification, likely because the VQ-VAE is a more complex model able to capture more subtle details. Nonetheless, the kNN tokens for classification work reasonably well, and may be a useful avenue to explore further in future work due to the relative ease of training the kNN in comparison to the VQ-VAE.
Following these results, all subsequent tests use a backbone with continuous inputs, ordering in the prediction head for pre-training and a classification loss using VQ-VAE tokens as targets in the pre-training. Unless explicitly stated the same backbone is used for all tests, this backbone is trained on the full JetClass training set using the VQ-VAE provided labels.
@@
no ordering continuous token classification $54.1\%$
order head continuous token classification $\mathbf{56.8\%}$
order backbone continuous token classification $53.4\%$
order head quantized token classification $51.1\%$
order head quantized kNN classification $49.3\%$
order head continuous kNN classification $56.2\%$

### v3

@@ I Introduction
While Artificial Intelligence (AI) and Machine Learning (ML) are already playing a major role in the analysis of high energy physics (HEP) data, the HEP community has yet to benefit from the self-supervised learning (SSL) based approaches to building large foundation models (FM) Bommasani et al. (2022) that have been pioneered in natural language processing (NLP) Lewis et al. (2019); Devlin et al. (2019); OpenAI (2023); Brown et al. (2020) and computer vision (CV) Dosovitskiy et al. (2021); Caron et al. (2021); Bao et al. (2022). FMs, as opposed to task specific ML models, are pre-trained in generic ways, such that they are useful for a range of downstream tasks. FMs often use SSL to pre-train models on vast data sets in order to learn generic representations of the data. Such models can then be efficiently fine-tuned with small datasets for a variety of downstream tasks. The self-supervised pre-training of a FM produces a model that is also referred to as the “backbone”, as it can serve as the information extraction component for downstream models. This concept significantly expands the possibilities for learning robust and meaningful data representations. The knowledge encoded in this way can be readily applied in downstream tasks which use this representation as input.
@@
The FM with a SSL pre-training approach offers important advantages for HEP. Unlike supervised learning, which typically acquires limited domain representations and focuses on a few key features for high prediction accuracy that must be learned anew for each task, SSL aims to learn generic representations summarizing domain features that prove useful across various downstream tasks. SSL tasks can be formulated on unlabeled data. In the HEP context, this may not only decrease the reliance on labeled simulated data sets but also potentially helps mitigate uncertainties related to domain shift when training models on imperfect simulations. However, this approach also has several major challenges for HEP. SSL strategies are data type specific, so new methods must be developed. These models also represent a scale in both model size and data size that have not been addressed in HEP. In this work, we aim to take the first steps towards building such a HEP foundation model, focusing on developing HEP data specific SSL strategies, understanding how these models perform when fine-tuned for various downstream tasks, and keeping an eye on how well such strategies may scale in the future. We propose a masked particle modeling (MPM) scheme, akin to masked language modeling (MLM) in NLP, for self-supervised learning on unlabeled data consisting of sets of particles in a collider physics environment. In doing so, we propose a novel scheme to apply masked modeling strategies to unordered sets of inputs.
[…]
@@
This work aims to generalize the language-inspired MLM-type training scheme to HEP scientific data, and to show that in the HEP context this can be an effective strategy for developing a useful FM capable of being fine-tuned for various down-stream tasks. The paradigm involves extracting semantic meaning and understanding of the whole by predicting the missing (masked) pieces, referred to as tokens, thereby considering the collective impact of individual input elements. Generic data sets, such as the ones in HEP, cannot be directly mapped to a sequence of pieces in analogy to words in a sentence as they are continuous and unordered, i.e. not a sequence. The challenges of using continuous elements, i.e. features of particles like momentum, instead of a discrete dictionary of words, and of operating on unordered sets of inputs will be addressed in the development of the MPM scheme. To do so, we examine MPM in collider physics data sets consisting of jets, or collimated streams of particles produced by high-energy quarks and gluons.
@@ II Related Work
Foundation models, such as Masked Language Models (BART Lewis et al. (2019), BERT Devlin et al. (2019)), Generative Pre-trained Transformer (GPT) OpenAI (2023); Brown et al. (2020), Vision Transformer (ViT) Dosovitskiy et al. (2021), DINO Caron et al. (2021) and their combinations, such as DALLE Ramesh et al. (2021), Flamingo Alayrac et al. (2022) and others have primarily been explored in the domains of language and vision. We refer readers to the recent review Bommasani et al. (2022) for an overview. Most closely related to this work is the BERT model Devlin et al. (2019), which uses the masking and prediction of missing words as a pre-training task, and the BEiT model Bao et al. (2022), which adapts the masked language modeling method to images by masking and predicting patches of input images. On masked modeling schemes for data which consists of unordered sets of inputs, the impact of removing positional information in masked image modeling was examined in Ref. Chen et al. (2021), and using position as a target when processing unordered image patches were explored in Ref. Zhai et al. (2022). The first steps in developing foundation models for science have been developed in e.g. protein biology Rives et al. (2021), molecular chemistry Ross et al. (2022); Pan (2023), and cosmology Lanusse et al. (2023); Walmsley et al. (2022), showing their ability to learn informative representations that are useful in these domains for various downstream tasks.
The first steps in self-supervised learning on jets was explored in Refs. Dillon et al. (2022, 2023); Tombs and Lester (2022), largely focusing on contrastive pre-training using augmentations of jets. Pre-training strategies through masking particle type information have also been explored in Ref. Kishimoto et al. (2023). Transformer models were trained on large jet datasets for classification in a supervised setting in Ref. Qu et al. (2022a); Mikuni and Canelli (2021) and several transformer-based applications have since been developed (for example, see Refs. K ach et al. (2022); Kansal et al. (2023); Fenton et al. (2022); ATLAS Collaboration (2023); Smith et al. (2023); Tomiya and Nagai (2023); K ach and Melzer-Pellmann (2023); Raine et al. (2023a)). Transformers have also been used for auto-regressive density estimation and jet classification Finke et al. (2023); Butter et al. (2023). Notably, Ref. Finke et al. (2023) also explored the discretization of continuous particle features to form jet sequences, which we examine in this work.
@@ III.1 Masked Particle Modeling
The representation of a jet as an unordered set of particles, each characterized by an ordered collection of continuous features, lends itself to interpretation within the masked training paradigm reminiscent of MLMs. In this analogy, each particle within the jet can be viewed as a representation akin to a token in a block of text, despite the notable difference that particles form unordered sets, while text is sequential.
@@
We can consider jets with at most $N$ constituent particles to be described as a set $X=\{x_{i}\}_{i=1}^{N}\in\mathcal{X}$ where each $x_{i}$ is some representation of a particle and $\mathcal{X}$ is the set of possible jets. A dataset is then a collection of $K$ jets $\{X^{j}\}_{j=1}^{K}$. The MPM objective can be phrased as partitioning the set of particles in each jet $j$ into a masked set $\mathcal{M}_{x}^{j}=\{x_{i}^{j}\}_{i\in\mathcal{M}^{j}}$ over indices $\mathcal{M}^{j}$ of masked elements, and an unmasked set $\mathcal{U}_{x}^{j}=\{x_{i}^{j}\}_{i\in\mathcal{U}^{j}}$ over indices of non-masked elements. A masking strategy is used for mapping $\mathcal{M}_{x}^{j}\rightarrow\mathcal{M}_{m}^{j}=\{m_{i}^{j}\}_{i\in\mathcal{M}^{j}}$ where each $m_{i}^{j}$ is a learnable vector. For convenience, we drop the jet index $j$ as the procedure is repeated for all jets. The goal is to define a parametric function $f_{\theta}:\mathcal{X}\to\mathbb{R}^{N\times d}$, which maps a jet to a $d$-dimensional representation for each of the $N$ particles in the jet, and a loss $\mathcal{L}$ such that minimizing $\mathbb{E}_{\mathcal{X}}[\mathcal{L}(\mathcal{M}_{x},f_{\theta}(\mathcal{M}_{m},\mathcal{U}_{x}))]$ results in a function $f_{\theta}$ that is useful for downstream tasks. As the goal of the pre-training is to recover information about each masked particle, we use a per-particle loss function of the form
@@
The particles that form a jet are permutation invariant, and so it is natural to use a backbone model that is permutation equivariant. However, if the same learnable vector is used to replace all of the particles in the masked set, then the output of the model will be equivalent at every masked position. In Fig. 1 this would mean that if $m_{1}=m_{2}$ then $h_{2}=h_{4}$, because permutation equivariant functions can not distinguish between two input particles with the same values. This exposes a redundancy that can only be removed by inducing an ordering, or by adding attributed connections between particles. The redundancy makes it impossible to achieve zero error on any single prediction unless the original value of every particle in the masked set is identical. However, the redundancy does not make MPM fruitless, as learning the density over the masked set of particles is still a useful and difficult task.
@@ III.2 Making Tokens From Particles
A scheme for defining context dependent tokens from continuous inputs was developed in the masked image modeling approach of the BEiT model Bao et al. (2022). In this case, labels of different image patches were defined using a Vector Quantized Variational AutoEncoder (VQ-VAE) Oord et al. (2017). A VQ-VAE uses an encoder to map a set of inputs to latent vectors, which are subsequently projected onto the nearest element within a finite codebook. These codebook vectors are then decoded back to the original inputs. The codebook vectors are trained simultaneously with the encoder and decoder. The use of transformer encoder and decoder ensures that information from all input elements are used to define the predictions. This process incorporates input-wide context into the codebook prediction for each element within the input set. In MPM, each particle is encoded to a single codebook element, where the encoding is performed conditionally on all other particles in the jet. The index of the codebook element to which each particle is mapped is then used as the target label during pre-training. The VQ-VAE model is only used during pretraining.
We also explore using the K-means algorithm MacQueen et al. (1967) to define labels for the pre-training. We use the K-means++ Arthur and Vassilvitskii (2007) algorithm as implemented in the scikit learn library Buitinck et al. (2013) to define the clusters. After training, each cluster is assigned an index, and the index is used as target labels. This approach is explored as a replacement for the VQ-VAE defined labels. Unlike the VQ-VAE labelling, the K-means approach is context independent. Further studies into the differences between the VQ-VAE and K-means would be beneficial.
@@ V Experiments
The same transformer encoder architecture is used for all backbone models. We used transformer blocks based on the Normformer Shleifer et al. (2021). Eight transformer blocks are used with a model dimension of $1024$, the input nodes are embedded linearly into a $512$ dimensional space, and the model contains a total of 40 million parameters. The pre-training head is a single transformer block followed by a softmax for classification or a linear layer for regression. The fine-tuning classification head is a weighted average over all output dimensions (i.e. a linear transform applied independently to each particle representation and then averaged) followed by another linear layer and softmax. Models are trained with AdamW Kingma and Ba (2014); Loshchilov and Hutter (2017) with a learning rate of $10^{-4}$ and weight decay of $0.01$. All pre-trained models are trained for five epochs on the full JetClass data set. For supervised training, all models are trained for up to $50$ epochs with early stopping and a patience of five epochs. The norm of the gradient vectors are clipped to five. The from scratch supervised models use the same architecture as the pre-trained models: eight transformer blocks followed by a weighted average and softmax for classification. The model architecture is similar to that used by ParT Qu et al. (2022a), which is state of the art. The ParT model uses more features than we use in this work, and so achieves higher accuracy. More details can be found in App. A.
@@
To define target labels and for input quantization, 512 possible token values are used. This corresponds to 512 vectors in the VQ-VAE codebook, or to 512 clusters for the K-means.
All the code used to run these experiments is publicly available at github.com/rodem-hep/mpm.
@@ V.1 Token Creation with a VQ-VAE
One challenge with training VQ-VAEs is the possibility of codebook collapse, where only a few of the codebook vectors are used to encode particle features. In the case of collapse to a single vector, the model would effectively be learning an average representation, which would not aid pre-training. To address this the prescriptions outlined in Ref. Huh et al. (2023) are applied as described in App. B. After training we find that the codebook has $\sim 80\%$ utilization. The quantized vectors, defined by their index or codebook vector, will be referred to as tokens.
@@ V.2 Discretization and Ordering
[…]
@@
The results of the comparison can be seen in Tab. 1 where five different random seeds were used during the fine tuning stage, but with only a single backbone trained in each case. The mean result is shown in the table, with a standard deviation of O(0.01%) for each result. A clear benefit to ordering only in the model head is observed over ordering the backbone inputs or not ordering anywhere in the model. In this case, ordering the prediction head enables the symmetry over masked elements to be broken and thus make more precise per-particle predictions rather than a joint posterior over all masked particles. Input quantization is observed to hurt model performance, owing to the loss of information when summarizing continuous features with discrete tokens. Using tokens, either from the VQ-VAE or the K-means, as classification targets is found to be substantially better than regressing particle features as a prediction task. This is likely due to the regression giving only a point prediction of average particle properties under the posterior, which provides little predictive power for posteriors with multiple modes. The classification loss is much more flexible and able to make use of the prediction of the full posterior over tokens. The VQ-VAE tokens outperformed the K-means tokens for classification, likely because the VQ-VAE is a more complex model able to capture more subtle details. Nonetheless, the K-means tokens for classification work reasonably well, and may be a useful avenue to explore further in future work due to the relative ease of training the K-means in comparison to the VQ-VAE.
[…]
@@
no ordering continuous VQ-VAE classification $54.1\%$
order head continuous VQ-VAE classification $\mathbf{56.8\%}$
order backbone continuous VQ-VAE classification $53.4\%$
order head quantized VQ-VAE classification $51.1\%$
order head quantized K-means classification $49.3\%$
order head continuous K-means classification $56.2\%$